# references.bib

### v1

@@ Acknowledgements
The work of G. Peyré was supported by the European Research Council (ERC project WOLF) and the French government under the management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PRAIRIE-PSAI). The work of F.-X. Vialard was supported by the Bézout Labex (New Monge Problems), funded by ANR, reference ANR-10-LABX-58.
@@ Appendix A Positive definite kernels and RKHS
$\forall\omega,\omega^{\prime}\in\Omega,\quad\kappa(\omega,\omega^{\prime})\coloneqq\int_{X}\phi(\omega,x)\phi(\omega^{\prime},x)\mathrm{d}\rho(x)=\langle\phi(\omega,.),\phi(\omega^{\prime},.)\rangle_{L^{2}(\rho)}.$
Thus, associated to $\kappa$ is a (uniquely defined) structure of Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$ with scalar product $\langle.,.\rangle_{\mathcal{H}}$, that is a Hilbert space of functions on $\Omega$ such that [steinwart2008support, Def. 4.18]: (i) $\kappa(\omega,.)\in\mathcal{H}$ for every $\omega\in\Omega$ and (ii) the following reproducing property holds:
$\forall h\in\mathcal{H},\omega\in\Omega,\quad h(\omega)=\langle h,\kappa(\omega,.)\rangle_{\mathcal{H}}.$
@@
$\forall h\in\mathcal{H},\quad\int_{\Omega}h\mathrm{d}\nu=\langle h,M_{\kappa}(\nu)\rangle_{\mathcal{H}}.$ (42)
Equivalently, the kernel embedding is given by the Böchner intgral $M_{\kappa}(\nu)=\int_{\Omega}\kappa(.,\omega)\mathrm{d}\nu(\omega)\in\mathcal{H}$. This embedding defines a discrepancy between measures by seeing them as element of the Hilbert space $\mathcal{H}$. For two measures $\nu,\nu^{\prime}\in\mathcal{M}(\Omega)$ the Maximum Mean Discrepancy (MMD) between $\nu$ and $\nu^{\prime}$ is defined as [muandet2017kernel, gretton2012kernel]:
$\mathrm{MMD}_{\kappa}(\nu,\nu^{\prime})=\left\|M_{\kappa}(\nu)-M_{\kappa}(\nu^{\prime})\right\|_{\mathcal{H}}.$
Alternatively, Assumption 3 is sufficient to ensure the convolution $\Phi\star:\mathcal{M}(\Omega)\to L^{2}(\rho)$ defined in eq. 7 is a bounded operator and by construction we have:
@@
The discrepancy $\mathrm{MMD}_{\kappa}$ is in particular a distance between measures whenever the kernel $\kappa$ is universal, that is when the associated RKHS $\mathcal{H}$ is dense in the space of continuous functions on $\Omega$ [sriperumbudur2011universality]. One can show this condition is equivalent to an injectivity assumption on the operator $\Phi\star$.
Lemma A.1.
@@
$0=\int_{\Omega}\left(\int_{\mathbb{R}^{d}}\phi(\omega,x)\alpha(x)\mathrm{d}\rho(x)\right)\mathrm{d}\nu(\omega)=\langle\alpha,\Phi\star\nu\rangle_{L^{2}(\rho)}.$
Hence $\Phi\star\nu=0$, implying $\nu=0$ and thus that =$\mathcal{H}^{\perp}=\{0\}$ i.e. $\mathcal{H}$ is dense in $\mathcal{C}^{0}(\Omega)$. For the converse implication, assume that $\mathcal{H}$ is dense in $\mathcal{C}^{0}(\Omega)$ and consider some $\nu\in\mathcal{M}(\Omega)$ s.t. $\Phi\star\nu=0$. Then for every $\alpha\in L^{2}(\rho)$ we have $\langle\alpha,\Phi\star\nu\rangle_{L^{2}(\rho)}=0$ and by similar calculations this implies $\nu\in\mathcal{H}^{\perp}$ i.e. $\nu=0$. ∎
@@
Then $J_{\mu}$ is compact, self-adjoint and positive operator and, by the spectral theorem, it can be diagonalized in an orthonormal basis $(e_{i})_{i\geq 0}$ of $L^{2}(\mu)$ with associated eigenvalues $(\lambda_{i})_{i\geq 0}$ s.t. $\lambda_{1}\geq\lambda_{2}\geq...\geq 0$. In particular, $J_{\mu}=\Phi_{\mu}^{*}\Phi_{\mu}$ with $\Phi_{\mu}:L^{2}(\mu)\to L^{2}(\rho)$ the feature operator defined in eq. 3, thus $(\sqrt{\lambda_{i}})_{i\geq 0}$ are the (right) singular values of $\Phi_{\mu}$ (which is a compact operator) and, if $\Phi\star$ is injective (i.e. $\mathcal{H}$ is universal according to Lemma A.1), then $\lambda_{i}>0$ for every $i\geq 0$. Mercer’s theorem gives a representation of the kernel $\kappa$ and of the associated RKHS $\mathcal{H}$ in terms of ths eigenvalue decomposition [steinwart2008support, Thm. 4.51].
Theorem A.1 (Mercer representation of RKHSs).
@@
is the RKHS associated to the kernel $\kappa$ when provided with scalar product $\langle.,.\rangle_{\mathcal{H}}$ defined for $f=\sum_{i\geq 0}a_{i}\sqrt{\lambda_{i}}$ and $g=\sum_{i\geq 0}b_{i}\sqrt{\lambda_{i}}$ by $\langle f,g\rangle_{\mathcal{H}}=\sum_{i\geq 0}a_{i}b_{i}$.

### v2

@@ Acknowledgements
The work of G. Peyré was supported by the European Research Council (ERC project WOLF) and the French government under the management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PRAIRIE-PSAI). The work of F.-X. Vialard was supported by the Bézout Labex (New Monge Problems), funded by ANR, reference ANR-10-LABX-58. This work was performed using HPC resources from GENCI-IDRIS (Grant 2025-AD011013400R3).
@@ Appendix A Positive definite kernels and RKHS
$\forall\omega,\omega^{\prime}\in\Omega,\quad\kappa(\omega,\omega^{\prime})\coloneqq\int_{X}\phi(\omega,x)\phi(\omega^{\prime},x)\mathrm{d}\rho(x)=\left<\phi(\omega,.),\phi(\omega^{\prime},.)\right>_{L^{2}(\rho)}.$
Thus, associated to $\kappa$ is a (uniquely defined) structure of Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$ with scalar product $_{\mathcal{H}}$, that is a Hilbert space of functions on $\Omega$ such that [steinwart2008support, Def. 4.18]: (i) $\kappa(\omega,.)\in\mathcal{H}$ for every $\omega\in\Omega$ and (ii) the following reproducing property holds:
$\forall h\in\mathcal{H},\omega\in\Omega,\quad h(\omega)=\left<h,\kappa(\omega,.)\right>_{\mathcal{H}}.$
@@
$\forall h\in\mathcal{H},\quad\int_{\Omega}h\mathrm{d}\nu=\left<h,M_{\kappa}(\nu)\right>_{\mathcal{H}}.$ (62)
Equivalently, the kernel embedding is given by the Bochner integral $M_{\kappa}(\nu)=\int_{\Omega}\kappa(.,\omega)\mathrm{d}\nu(\omega)\in\mathcal{H}$. This embedding defines a discrepancy between measures by seeing them as element of the Hilbert space $\mathcal{H}$. For two measures $\nu,\nu^{\prime}\in\mathcal{M}(\Omega)$ the Maximum Mean Discrepancy (MMD) between $\nu$ and $\nu^{\prime}$ is defined as [muandet2017kernel, gretton2012kernel]:
$\mathrm{MMD}_{\kappa}(\nu,\nu^{\prime})\coloneqq\left\|M_{\kappa}(\nu)-M_{\kappa}(\nu^{\prime})\right\|_{\mathcal{H}}.$
Alternatively, Assumption 3 is sufficient to ensure the operator $\Phi\star:\mathcal{M}(\Omega)\to L^{2}(\rho)$ defined in Eq. 4 is bounded and by construction we have:
@@
The discrepancy $\mathrm{MMD}_{\kappa}$ is in particular a distance between measures whenever the kernel $\kappa$ is universal, that is when the associated RKHS $\mathcal{H}$ is dense in the space of continuous functions on $\Omega$ [micchelli2006universal, sriperumbudur2011universality]. One can show this condition is equivalent to an injectivity assumption on $\Phi\star$.
Lemma A.1 (see also [micchelli2006universal, Prop. 1]).
@@
$0=\int_{\Omega}\left(\int_{\mathbb{R}^{d}}\phi(\omega,x)\alpha(x)\mathrm{d}\rho(x)\right)\mathrm{d}\nu(\omega)=\left<\alpha,\Phi\star\nu\right>_{L^{2}(\rho)}.$
Hence $\Phi\star\nu=0$, implying $\nu=0$ and thus that =$\mathcal{H}^{\perp}=\{0\}$ i.e. $\mathcal{H}$ is dense in $\mathcal{C}^{0}(\Omega)$ by Hahn-Banach theorem. For the converse implication, assume that $\mathcal{H}$ is dense in $\mathcal{C}^{0}(\Omega)$ and consider some $\nu\in\mathcal{M}(\Omega)$ s.t. $\Phi\star\nu=0$. Then for every $\alpha\in L^{2}(\rho)$ we have $\alpha,\Phi\star\nu\right>_{L^{2}(\rho)}=0$ and by similar calculations this implies $\nu\in\mathcal{H}^{\perp}$ i.e. $\nu=0$. ∎
@@
Then $J_{\mu}$ is a compact, self-adjoint and positive operator and, by the spectral theorem, it can be diagonalized in an orthonormal basis $(e_{i})_{i\geq 0}$ of $L^{2}(\mu)$ with associated eigenvalues $(\lambda_{i})_{i\geq 0}$ s.t. $\lambda_{1}\geq\lambda_{2}\geq...\geq 0$. In particular, $J_{\mu}=\Phi_{\mu}^{\top}\Phi_{\mu}$ with $\Phi_{\mu}:L^{2}(\mu)\to L^{2}(\rho)$ the feature operator defined in Eq. 11, thus $(\sqrt{\lambda_{i}})_{i\geq 0}$ are the (right) singular values of $\Phi_{\mu}$ (which is a compact operator) and, if $\Phi\star$ is injective, then $\lambda_{i}>0$ for every $i\geq 0$. Mercer’s theorem gives a representation of the kernel $\kappa$ and of the associated RKHS $\mathcal{H}$ in terms of this eigenvalue decomposition [steinwart2008support, Thm. 4.51].
Theorem 6 (Mercer representation of RKHSs).
@@
is the RKHS associated to the kernel $\kappa$ when provided with the scalar product $_{\mathcal{H}}$ defined for $f=\sum_{i\geq 0}a_{i}\sqrt{\lambda_{i}}$ and $g=\sum_{i\geq 0}b_{i}\sqrt{\lambda_{i}}$ by $f,g\right>_{\mathcal{H}}=\sum_{i\geq 0}a_{i}b_{i}$.
Appendix B Radial basis function neural network on the $2$-dimensional torus
We performed numerical experiments to test the performance of the VarPro algorithm for the training of Radial Basis Function (RBF) neural networks. Notably, due to the particular structure of the architecture, the learning problem corresponds to performing a deconvolution, which has important applications in signal processing [de2012exact, duval2015exact].
The feature space is here the $2$-dimensional torus $\Omega=\nicefrac{{\mathbb{R}^{2}}}{{4\mathbb{Z}^{2}}}\subset\mathbb{R}^{2}$ and the data dimension is $d=2$. The RBF neural network architecture performs the convolution with a kernel $k:\mathbb{R}^{2}\to\mathbb{R}$ and corresponds to considering the feature map $\phi:(\omega,x)\mapsto k(\omega-x)$. We will use here the Laplace kernel $k:x\in\mathbb{R}^{2}\mapsto 8\exp(-\frac{1}{2}\|[x]\|)$, where $[x]$ represents the projection of $x$ in $\Omega=\nicefrac{{\mathbb{R}^{2}}}{{4\mathbb{Z}^{2}}}$. For features $\{\omega_{i}\}_{i=1}^{M}\in(\Omega)^{M}$ and outer weights $\{u_{i}\}_{i=1}\in\mathbb{R}^{M}$ the RBF neural network model reads:
$F_{\{(\omega_{i},u_{i})\}}:x\in\mathbb{R}^{2}\mapsto\frac{1}{M}\sum_{i=1}^{M}u_{i}k(\omega_{i}-x)=(k\star\hat{\nu})(x),$ (64)
where $\hat{\nu}=\frac{1}{M}\sum_{i=1}^{M}u_{i}\delta_{\omega_{i}}\in\mathcal{M}(\Omega)$ and $\star$ is the convolution operator. We consider a teacher feature distribution $\bar{\mu}_{\gamma}=\frac{1}{\bar{M}}\sum_{i=1}^{\bar{M}}\delta_{\bar{\omega}_{i}}$ for features $\{\bar{\omega}_{i}\}_{1\leq i\leq\bar{M}}\in\Omega^{\bar{M}}$ and the target signal $Y$ is thus:
$\forall x\in\mathbb{R}^{2},\quad Y(x)=\frac{1}{\bar{M}}\sum_{i=1}^{\bar{M}}k(\omega_{i}-x)=(k\star\bar{\mu}_{\gamma}).$
The teacher features are i.i.d. with $\bar{\omega}_{i}\sim\mu_{\gamma}=(\frac{1}{2}\delta_{\omega^{*}_{1}}+\frac{1}{2}\delta_{\omega^{*}_{2}})\star\pi_{\gamma}$, where $\omega_{1}^{*}=(-1,0)$, $\omega^{*}_{2}=(1,1)$ are two target modes and $\pi_{\gamma}$ is the product measure with density:
$\forall(z_{1},z_{2})\in\nicefrac{{\mathbb{R}^{2}}}{{4\mathbb{Z}^{2}}},\quad\pi_{\gamma}(z_{1},z_{2})\propto\frac{1}{1+\gamma\sin^{2}(z_{1}\pi/4)}\times\frac{1}{1+\gamma\sin^{2}(z_{2}\pi/4)}\,.$
The parameter $\gamma$ controls the shape of the distribution $\mu_{\gamma}$ such that in the large $\gamma$ limit one recovers $\mu_{\gamma}\simeq\mu_{\infty}\coloneqq\frac{1}{2}\delta_{\omega^{*}_{1}}+\frac{1}{2}\delta_{\omega^{*}_{2}}$. A scatter plot of the teacher measure $\bar{\mu}_{\gamma}$ and of the resulting teacher signal for $\gamma=100$ are shown in Fig. 10. Finally, we consider the input data $x$ to be distributed according to an empirical distribution $\hat{\rho}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ with i.i.d. standard gaussian samples $x_{i}\sim\mathcal{N}(0,\mathrm{Id})$. In this setting we consider training the model by performing a VarPro algorithm i.e. gradient descent over the reduced risk, as in Eq. 52.
Figure 10: Left: scatter plot the empirical teacher distribution $\bar{\mu}_{\gamma}$ for $\gamma=100$. Right: corresponding target signal.
Experimental setting
We test the performances of the VarPro algorithm (Eq. 52) for the training of a RBF neural network (Eq. 64) of varying width $M\in\{32,128,512,1024\}$. We use either the “biased” quadratic regularization $f_{b}:t\mapsto\frac{1}{2}t^{2}$ or the “unbiased” quadratic regularization $f_{u}:t\mapsto\frac{1}{2}|t-1|^{2}$ and we consider varying regularization strength $\lambda\in\{10^{-1},10^{-2},10^{-3}\}$. We consider different teacher distributions $\bar{\mu}_{\gamma}$ by changing the parameter $\gamma\in\{100,+\infty\}$. As in Section 6.1, we consider a number of data samples $N=4096\gg M$, a teacher of width $\bar{M}=4096\gg M$ (s.t. the approximation $\bar{\mu}_{\gamma}\simeq\mu_{\gamma}$ holds) and a stepsize $\tau=2^{-10}$.
Student of varying width
As before, we first investigate the role played by the width $M$ of the student in the training dynamic. For this purpose, we fix the regularization strength to $\lambda=10^{-3}$ and consider training RBF neural networks (Eq. 64) of varying width $M\in\{32,128,512,1024\}$ with the teacher distribution $\bar{\mu}_{\gamma}$, $\gamma=100$.
Fig. 11 reports evolution of the biased and unbiased reduce risk during training and Fig. 12 reports evolution of the distance to the teacher distribution $\bar{\mu}_{\gamma}$. As for our $1$-dimensional experiments, one can observe that the VarPro algorithm converges to lower values of the unbiased reduced risk when the width of the student increases. In turn, at convergence, this corresponds to learned feature distributions that approximate the teacher distribution with different levels of discretization. On the contrary, using the biased regularization $f_{b}:t\mapsto\frac{1}{2}t^{2}$ introduces a bias in the learned distribution.
Figure 11: Evolution of the reduced risk along iterations of gradient descent for a RBF neural network (Eq. 64) of width $M\in\{32,128,512,1024\}$. The regularization strength is $\lambda=10^{-3}$ and the regularization function is either $f_{b}:t\mapsto\frac{1}{2}t^{2}$ (left) or $f_{u}:t\mapsto\frac{1}{2}|t-1|^{2}$ (right). Plots are averages over $6$ independent runs.
Figure 12: Evolution of the MMD distance to the teacher distribution $\bar{\mu}_{\gamma}$ along gradient descent over the reduced risk for a RBF neural network (Eq. 64) of width $M\in\{32,128,512,1024\}$. The regularization strength is $\lambda=10^{-3}$ and the regularization function is either $f_{b}:t\mapsto\frac{1}{2}t^{2}$ (dashed) or $f_{u}:t\mapsto\frac{1}{2}|t-1|^{2}$ (plain). Plots are averages over $6$ independent runs.
Role of the regularization strength $\lambda$
We now investigate the role of the regularization strength $\lambda>0$. We thus consider training RBF neural networks (Eq. 64) of fixed width $M=1024$ with the teacher distribution $\bar{\mu}_{\gamma}$, $\gamma=100$, and we perform gradient descent over the reduced risk (Eq. 52) with varying regularization $\lambda\in\{10^{-1},10^{-2},10^{-3}\}$. Note that here, compared to our $1$-dimensional, the case of regularization lower than $\lambda=10^{-3}$ is numerically impracticable, at least with our choice of stepsize $\tau=2^{-10}$.
Evolution of the distance to the teacher distribution $\bar{\mu}_{\gamma}$ along training is reported in Fig. 13. As in our $1$-dimensional experiments, one can observe that the convergence speed gets slower when the regularization strength increases. There is also a significant change of behavior between $\lambda=10^{-1}$ and $\lambda\in\{10^{-2},10^{-3}\}$. In the former case the convergence seems to exhibit an algebraic rate, supporting the conclusions of Theorem 4, while in the latter the convergence rate is linear, indicating a behavior closer to the ultra-fast diffusion limit (Theorem 3). As for the $1$-dimensional case, this can be explained by the fact that $\lambda=10^{-1}$ is the order of magnitude of the most significant eigenvalues of the tangent kernel $K_{\mu}$ in Eq. 32. Thus, for higher values of $\lambda$, one enters in a high regularization regime where the reduced risk receive more influence from the MMD distance term than from the $f$-divergence term in Eq. 26. One also observes that the bias introduced in the case of the regularization $f_{b}:t\mapsto\frac{1}{2}t^{2}$ vanishes with the regularization strength $\lambda$, supporting the conclusions of our Proposition 3.1.
Figure 13: Evolution of the MMD distance to the teacher distribution $\bar{\mu}_{\gamma}$ along gradient descent over the reduced risk (Eq. 52) for a RBF neural network (Eq. 64) of width $M=1024$ with regularization $\lambda\in\{10^{-1},10^{-2},10^{-3},10^{-4}\}$. The regularization function is either $f_{b}:t\mapsto\frac{1}{2}t^{2}$ (dashed) or $f_{u}:t\mapsto\frac{1}{2}|t-1|^{2}$ (plain). Plots are averages over $6$ independent runs.
Role of the shape of the teacher distribution
Finally, we investigate the impact of the shape of the teacher distribution on the VarPro dynamic. We are particularly interested in the limit $\gamma=+\infty$ in which the teacher distribution is given by $\bar{\mu}_{\gamma}=\frac{1}{2}\delta_{\omega^{*}_{1}}+\frac{1}{2}\delta_{\omega^{*}_{2}}$. While such setting is not covered by our theory (in particular the ultra-fast diffusion equation is not necessarily well-posed), it is of interest to see if the VarPro algorithm is able to recover sparse feature representations.
In this context, we consider teacher distributions $\bar{\mu}_{\gamma}$ for $\gamma\in\{100,+\infty\}$ and train RBF neural networks (Eq. 64) of fixed width $M=1024$ with gradient descent over the reduced risk (Eq. 52) with the unbiased regularization $f_{u}:t\mapsto\frac{1}{2}|t-1|^{2}$ and $\lambda=10^{-3}$. Plots of the evolution of the reduced risk and of the MMD distance to the teacher distribution are reported in Fig. 14. As in our $1$-dimensional experiments, one can see that the convergence speed of VarPro deteriorates when $\gamma$ increases, both in terms of convergence of the risk and in terms of convergence of the learned feature distribution to the teacher’s. In case of a sparse teacher distribution ($\gamma=+\infty$), convergence towards the teacher seems to not necessarily be governed by a linear rate. Indeed, as the teacher distribution is not absolutely continuous, one could expect the comparison with ultra-fast diffusion dynamics to no longer hold. However, Theorem 4 could still apply, leading to an algebraic convergence rate.
Figure 14: Gradient descent over the reduced risk (Eq. 52) for a RBF neural network (Eq. 64) of width $M=1024$ with unbiased regularization $f_{u}:t\mapsto\frac{1}{2}|t-1|^{2}$, $\lambda=10^{-3}$ and teacher distributions $\bar{\mu}_{\gamma}$ for $\gamma\in\{100,+\infty\}$. Left: Evolution of the unbiased reduced risk. Right: Evolution of the MMD distance to the teacher distribution $\bar{\mu}_{\gamma}$. Plots are averages over $6$ independent runs.